Station SACV, Santiago Island, Cape Verde

SACV operational status:

2022.04.01

IDA engineers visited the station and restored communications between the digitizers and station computers on day 2022,085 (3/26).  They replaced the Advantech switch, cat-5 cables, Moxa hub, and Axiomatic power supply.  They also repaired a bad cell for the backup battery stack, upgraded the Q330 firmware, and established a cell link for the station.

2022.03.17

The station is down again due to communications issues between the digitizers and station computers (last data 2022,059-18:57).  IDA engineers will be visiting the station soon.

2021.12.01

Station operators visited the station and replaced both Q330s and fixed the cat-5 uplink cable. Reported that one of the Absolyte batteries is leaking.  IDA engineers are preparing for a possible trip in early 2022.

2021.08.26

The station is back online after the station operator visited the station and restored communications between the digitizers and station computers.

2021.07.27

The station is down due to communications issues between the digitizers and station computers.

2019.10.25

The station was down day 2019,295 to 298 due to issues with connections to the digitizers and the station computers.   IDA engineers worked with the station operator to resolve these issues.

2017.11.10

IDA engineers visited the station and replaced the KS54000 with a Trillium 360 seismometer, and installed a Hyperion Infrasound sensor, new station computers, and a weather transmitter.

2015.09.09

All channels from the KS54000 seismometer are non-seismic due to a faulty cable to the sensor.

2014.08.22

There was an extensive telemetry outage to the station from day 2014,108 to day 2014,234.  The station operator had suspected a broken phone cable to the station caused by a storm with heavy winds. Telemetry to the station was restored with no data loss after an IDA engineer visited the station and replaced a faulty power converter/data transfer box causing a disconnect between the telemetry and DSL modem.  While onsite the engineer also replaced the station computers and reconfigured the Setra barometer to input to the Q330-10 DAS.  The East-West channels from the KS54000 seismometer are still corrupted due to cross-talk between the two horizontal elements of the seismometer.

2013.06.03

Telemetry to the station has been problematic since April 2013.  There were several telemetry outages due to a faulty CVTELCOM cable to the station.  Once the cable was repaired telemetry was restored with no data loss.

2013.01.14

IDA station SACV has been upgraded to a Quanterra Q330HR system. Sensors installed include a Trillium 120 posthole seismometer, an Episensor strong motion sensor, and a Setra microbarograph. Upon arrival an IDA engineer noted that the station not acquiring data since it went down day 2012,208, and data from the KS54000 seismometer was non-seismic once the station was brought back online.  The cable from the KS54000 sensor to the KS-term box had corroded pins on the plug.  The engineer soldered the cable to the KS-term box which restored data acquisition for the KS54000.  However, the East-West component has been corrupted due to cross-talk between the two horizontal elements of the KS-54000 seismometer.

2012.07.30

IDA station SACV is down as of day 2012,208-20:41 due to power problems. Replacement batteries for the power system are being ordered to ship to the station.

2012.05.21

The voltage output of all three channels of the KS-54000 seismometer fell by a factor of 2.0 on 2011,167-15:00, after the DAS was replaced with one that contains an RT593 circuit board. The previous DAS contained 3 RT303 circuit boards with a least count of twice that of a DAS with an RT593 circuit board.

2011.06.17

The station is back on-line as of day 2011,167-14:46 after an IDA engineer visited the station and performed maintenance on the power system. Telemetry was restored once an internet outage on the island ceased.

2011.03.18

The station operator reports going to the station and determining that there is a power problem. IDA engineers are still working with the station operator to resolve the problems at the station.

2011.03.02

The station operator reports replacing the DAS on day 2011,061. Telemetry to the station remains down. IDA engineers are currently working with the station operator to restore telemetry so the status of the station can be determined.

2010.08.24

The station is experiencing problems with the DAS that are causing excessive gapping in the data since day 2010,235.

2009.12.09

The station is back on-line as of day 2009,343 after an IDA engineer visited the station and upgraded the power system.

2009.11.25

The station is down as of day 2009,329-23:36. Power problems are suspected and an IDA engineer is enroute to the station to upgrade the power system.

2009.08.05

The station was down from day 2009,209-05:05 to 2009,217-12:38 due to problems with the power supply. The station is currently up, but the problems with power are suspected to persist.

2009.07.10

The station operator visited the station and reported the station is operating normally. Telemetry to the station is down, and the station operator is contacting the service provider.

2009.06.29

The station is down as of about day 2009,180 for reasons yet to be determined. A power outage at the station is suspected.

2008.11.24

The station is up as of about day 2008,329 after the station operator replaced the GPS clock.

2008.10.20

The station is down due to power problems, and problems with the GPS clock.

2008.10.06

The station is currently up. However, there are still problems with power to the station.

2008.09.29

The station is down due to persistant power problems.

2008.09.17

The station is back on-line as of about day 2008,257 after the station operator restored communication with the station computer.

2008.09.09

The station is down due to problems with the power system.

2008.09.02

The station is back online as of about day 2008,244 after an IDA engineer visited the station and fixed problems with the ISP. However, persistant power problems remain at the station.

2008.08.18

The station is down as of about day 2008,203 due to problems with the ISP. A replacement ISP has been shipped to the station from UCSD.

2008.06.02

The station is back online after IDA engineers visited the station and fixed the power system, installed an new ISP, and restored telemetry to the station.

2007.06.30

An IDA engineer visited the station and retored power and investigated the status of equipment. There is a faulty ISP that will be replaced soon to bring the station back online.

2005.02.28

Station is down due to problems with power to the station. The solar batteries are depleted.

2004.08.09

Station is up (as of July 17) after IDA engineer replaced faulty XCOMM and RT303 boards in the DAS, and modems. Telemetry restored after station operator rebooted the ISP.

2004.06.08

Station is down due to communication problems between the DAS and ISP.

2004.05.06

GPS clock works normally since 2004:127-11:30.

2004.04.27

Telemetry data showed GPS clock problem since 2004:118-17:21.

2004.03.11

Station is back up after an IDA engineer reconfigured the connection between the DAS, station processor, and GPS clock.

2003.12.01

Station is down (11/27) due to problems with the DAS and external clock. There are frequent data gaps due to a RT303 board failure for the BBZ channels. After the DAS was rebooted on 2003:331-17:09 (11/27) there were numerous error messages and then communication to the DAS was lost.

2003.11.24

Telemetry link restored on 11/17. NRTS data available as of 2003:321-17:59 (11/17).

2003.09.08

Telemetry down as of 2003:231-06:24 (Aug. 19), however, station remains operational.

2003.05.27

Telemetry restored during visit by an IDA engineer. Problem resolved after a cross cable was connected to the radio at the gateway by station operators. During visit an IDA engineer restored the gateway hard drive, upgraded DAS firmware to DSPV777, and replaced the digitizer’s RT303 card.

2002.09.04

Recording problem corrected by removing faulty RT303 card. Station operational with two 24 bit channels. Telemetry problems persist. EPROMS replaced 2002.09.03.

2002.08.20

Station is down due to problems with the DAS.

2002.04.22

Station is operating normally, but the telemetry circuit is down. New radio hardware at the repeater site is now in place but must be reconfigured by the operator.